OpenStudy (anonymous):

Using half-angle identities, determine: 1. cos(x/2) if tanx = 12/5 and pi < x < 3pi/2 2. in which quadrant

OpenStudy (anonymous):

tanx = 12/5 p^2 + b^2 = h^2 h= 13 cosx = 5/13 x = (inverse) cos5/13...

OpenStudy (anonymous):

it is 67.38 the x

OpenStudy (anonymous):

but in third quad cos is negative

OpenStudy (anonymous):

so cosx = -5/13

OpenStudy (anonymous):

and third quadrant

OpenStudy (anonymous):

thanks for the answer but im looking for the exact value of cos(x/2). I think that using half-angle identities is necessary

OpenStudy (anonymous):

i forgot it is cosx/2 sorry let me do this again

OpenStudy (anonymous):

cosx = -5/13 cos 2* x/2 = 2cos^2 x/2 -1 cosx +1 = 2cos^2 x/2 1 - 5/13 = 2 cos^2 x/2 sqrt(8/(13 *2)) = cosx/2 hers is your answer and pi<x<3pi/2\ hence pi/2<x < 3pi/4

OpenStudy (anonymous):

which is second quadrant

OpenStudy (anonymous):

whats the basis for finding the quadrant?

OpenStudy (anonymous):

if it is in quadrant 2, then why is tanx = 12/5 and not negative?